A shoe store sold 340 pairs of shoes in one day. The number of pairs of athletic shoes sold was 4 more than twice the number of pairs of dress shoes sold. How many pairs of athletic shoes were sold?
Answer by stanbon(75887)

A shoe store sold 340 pairs of shoes in one day. The number of pairs of athletic shoes sold was 4 more than twice the number of pairs of dress shoes sold. How many pairs of athletic shoes were sold?
--------------------------
Equations:
a + d = 340
a = 2d+4
------------------------
Substitute for "d" and solve for "a":
d = 340-a
-----
a = 2(340-a) + 4
a = 684 - 2a
3a = 684
a = 228 (# of athletic shoes sold)
=========================================
Cheers,
Stan H.